OpenStudy (anonymous):

Find the amplitude y = -5 sin1/2x.

OpenStudy (anonymous):

so is the answer 5pi/2 ?

OpenStudy (ankit042):

value of sin function lies between +1 and -1 now try to answer

OpenStudy (jdoe0001):

from the picture, what do you think will be the amplitude for \(\bf -5sin\left(\frac{1}{2}x\right)\) ?

OpenStudy (anonymous):

I think I just confused myself

OpenStudy (jdoe0001):

@maywadee313 sine and cosine functions, no matter what angle you throw at them, they will always return a value that's, as ankit042 said, -1 or 1, check your Unit Circle or even your calculator for that

OpenStudy (jdoe0001):

-1 and between and 1 that is

OpenStudy (jdoe0001):

never higher, never lower

OpenStudy (anonymous):

oh alrighty

OpenStudy (jdoe0001):

so if that's the case, what do you think "y" is when \(\bf y = -5sin\left(\frac{1}{2}x\right)\)

OpenStudy (jdoe0001):

keep in mind that -5 will multiply against a number, that number will "never be higher than 1, lower than -1" so what do you think will be the HIGHEST value "y" will achieve? so what do you think will be the LOWEST value "y" will achieve?

OpenStudy (anonymous):

so is the answer pi/5?

OpenStudy (ankit042):

NO how did you get pi/5..can you explain?

OpenStudy (ankit042):

read the explanation by @jdoe0001

OpenStudy (jdoe0001):